3,561 views
1 1 vote
a fair coin is tossed 10 times ,wt is the probability that head will come up in first two toss of coin??

2 Answers

Best answer
2 2 votes

"a fair coin is tossed 10 times ,wt is the probability that head will come up in first two toss of coin"

We do not even need to consider remaining tosses.

So Total Set of outcomes of First 2 tosses is = { HH,HT,TH,TT}

Required outcome = {HH}

Answer -> cardinality {HH} / cardinality {HH,HT,TH,TT}  = 1/4 = 0.25

• selected by
5 5 votes
No of way(s) to fillFirst two places = 1

Rest tosses may result Head or Tail = 2*2*2*.......2 = 2^8 = 256

Total sample space = 2*2*2.....*2 = 2^10 = 1024

P (1st two tossesresult Head) = 256/1024 = 1/4 = 0.25
Position:
Show:

Related questions

32 32 votes
3 answers 3 answers
15.6k
15.6k views
gatecse asked Sep 21, 2014
15,565 views
Let $f(x)$ be the continuous probability density function of a random variable $x$, the probability that $a < x \leq b$, is :$f(b-a)$$f(b) - f(a)$$\int\limits_a^b f(x) dx...
61 61 votes
5 answers 5 answers
14.9k
14.9k views
Kathleen asked Sep 17, 2014
14,875 views
A program consists of two modules executed sequentially. Let $f_1(t)$ and $f_2(t)$ respectively denote the probability density functions of time taken to execute the two ...
13 13 votes
1 1 answer
1.3k
1.3k views
GO Classes asked May 5
1,290 views
A machine produces a random score $T$ between $0$ and $2$ with probability density function $$f(t)= \begin{cases}\frac{3t^2}{8}, & \text { if } 0<t<2 \\\\ 0, & \text { ot...
7 7 votes
1 1 answer
893
893 views
GO Classes asked May 5
893 views
Let $Y$ be a continuous random variable with density,$$f_Y(y)= \begin{cases}2y, & \text { if } 0<y<1 \\\\ 0, & \text { otherwise }\end{cases}$$Given $Y=y$, the conditiona...